Find locus of the centres of the circle such that the point (2,3) is the mid point of the chord 5x+2y=16.

Grade:12th pass

1 Answers

Arun
6 years ago
Let locus lie on
Y = mx +C
Slope of a line = -5/2
Hence slope of its perpendicular line = 2/5
Hence
y = 2x/5  +C
5y = 2x + 5C
(2, 3) satisfies this
C = 11/5
Hence
Line will be
5y = 2x + 5*11/5
2x - 5y + 11 = 0